OpenStudy (anonymous):

how do you evaluate C(9,7)

OpenStudy (amistre64):

P(9,7)/7!

OpenStudy (amistre64):

9.8.7.6.5.4.3 ------------ 7.6.5.4.3.2.1 simplify to your hearts desire

OpenStudy (anonymous):

C(n,k) can also be written: \[C(n,k) = \frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

What's The Answer ? A) 181,440 B)72 C)36 D)63

OpenStudy (amistre64):

please dont ask for an answer that you can work out for yourself ...

OpenStudy (amistre64):

\[C(n,k) = \frac{n!}{k!(n-k)!}\] \[C(n,k) = \frac{n!/(n-k)!}{k!}\] \[C(n,k) = \frac{P(n,k)}{k!}\]

OpenStudy (anonymous):

\[C(9,7) = \frac{9!}{7!(9-7)!} = \frac{9!}{7!2!} = \frac{7!*8*9}{7!2!} = \frac{8*9}{2}\]

OpenStudy (anonymous):

Thanks (: really Appreciate it..